A decapeptide (Mol. Wt. 796) on complete hydrolysis gives glycine (Mol. Wt. 75), alanine and phenylalanine. Glycine contributes 47.0% to the total weight of the hydrolysed products. The number of glycine units present in the decapeptide is  
In this question how would we know that 9 molecules of water had to be add ?
 
 

Grade:11

1 Answers

sohail
5 years ago
47 % of (796+162)=958*47/100=450approx so no of glycene units = 450/75=6 approx here 162 is taken for the weight of 9 water molecules required for breaking the decapeptides